f(x)=2x+4. Find f(-3) and f(5).
Answer by stanbon(75887)

f(x)=2x+4. Find f(-3) and f(5).
-----
f(-3) = 2(-3)+4 = -6+4 = -2
------------------------------
f(5) = 2(5)+4 = 14
=======================
Cheers,
Stan H.
===========